The length of a rectangular box is 1 inch more than twice the height of the box, and the width is 3 inches more than the height. If the volume of the box is 126 cubic inches, find the dimensions of the box. L=2x+1; H=x; W=x+3, V=126 cu. in.
Answer by scott8148(5891)

(2x+1)(x)(x+3)=126 __ 2x^3+7x^2+3x-126=0

factors of 126 are 7, 3, 3, 2, 1

factors of 2 are 2, 1

testing factors shows 3 as a root of the equation, so x-3 is a factor

(x-3)(2x^2+13x+42)=0 __ the discriminant of 2x^2+13x+42 is negative, so the other two roots are not real

so, x=3 and the dimensions are L=7, H=3, W=6